OpenStudy (anonymous):

V varies inversely with T and V = 12 when T = 6. Which equation shows this relationship?

OpenStudy (whpalmer4):

\[V = \frac{k}{T}\]\[12=\frac{k}{6}\]Solve for \(k\)

OpenStudy (whpalmer4):

\[V=2T\] is direct variation

OpenStudy (anonymous):

So thats the answer?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Look at my first post. Can you solve for \(k\)?

OpenStudy (anonymous):

k=24

OpenStudy (whpalmer4):

Show me your work...

OpenStudy (anonymous):

OH SORRY 72

OpenStudy (whpalmer4):

Yeah, much better! You should check it yourself..."does 24/6 = 12? Uh, nope, self, let's see where we went wrong!" :-)

OpenStudy (anonymous):

>.< Yea stupid me :P

OpenStudy (anonymous):

V=72/T?

OpenStudy (anonymous):

@shamil98

OpenStudy (whpalmer4):

Let's test it out: \[V=\frac{72}{T}\]\[V=12,\,T=6\]\[12=\frac{72}{6} = 12\checkmark\]

OpenStudy (anonymous):

So it is right!

OpenStudy (whpalmer4):

I encourage you to test your answers to the extent possible...

OpenStudy (anonymous):

Thanks so much !!!!!

OpenStudy (anonymous):

Yea sorry :( not good at these